NDOS CommissionMtng 20120507

May 7 Commisioning Meeting: Jan, Nick, Susan, Minerba, Jon Leon, Andrew Zukai, Nate, Athans David Peter, Mat

1.  Beam and Detector Update
Production DCMs and TDUs are installed on the detector
Single primary timing chain master->slave->slave with 2 branches on each slave
Secondary Hardware in place. Cabling later

2.  DataCheck Update Jan
Jan report that t00 (Numi) is updating (Is it an A9 or MIBs7)
Sunday afternoon, lost channels do to t-storm.
Email notification every 5 min.

3.  Installation Plans for the week
APD and FEB installations on DiB 3

4.  DAQ Update and Plans for the week
No Updates
Partitioning tests.

5.  DCS Update and Plans for the week (Athans)
Email notification soon.
Writing alarms into the database (in 3 weeks)

6.  AOB
New shifter procedure for the remote skeleton shift